Q.2 The Kenneth State organization sells a product for which the annual demand is 5000 units.

The average holding cost is $ 1.00 per unit per year, and the ordering cost is $ 100 per order.

  1. What is the economic order quantity for the product?
  2. What is the total annual cost for this EOQ?

Q.2 The Kenneth State organization sells a product for which the annual demand is 5000 units.

The average holding cost is $ 1.00 per unit per year, and the ordering cost is $ 100 per order.

  1. What is the economic order quantity for the product?
  2. What is the total annual cost for this EOQ?

***Please do NOT use handwriting***
Answer #1

a) The economic order quantity (EOQ) is

D = Annual demand = 5000

S = Ordering cost per unit = $100

H = total holding cost = $ 1

= 1000 units

b) The total annual cost for this EOQ C = (EOQ/2)*H + (D/Q)*S = (1000/2)*1 + (5000/1000)*100 = $ 1000.